# Online Smoothed Demand Management

### v1

@@ 4 A Unified Competitive Algorithm
4.1 The PAAD Algorithm
Figure 2: Diagram of the intuition behind the PAAD algorithm. Each driver accounts for a single unit of base or flexible demand, making local decisions based on its own state and role that are aggregated into a single global decision. Purchasing decisions $x_{t}$ govern how much to purchase at the current market price, and delivery decisions $z_{t}$ dictate how much demand is satisfied at the current time (subject to feasibility constraints).
@@ 4.1 The PAAD Algorithm
Given a base driver with size denoted by $d$, the threshold function $\phi_{b}$ for OSDM-S is defined as follows, where $\alpha$ is the competitive ratio (see Thm. 4.1):
@@
$\psi(v)=p_{\max}(c+\varepsilon)+2\delta+\Big(\frac{p_{\max}(c+\varepsilon)+2\delta}{\alpha^{\prime}}-\Big(p_{\max}(c+\varepsilon)+\frac{2\delta}{T}\omega\Big)\Big)\exp\left(\frac{v}{\alpha^{\prime}d}\right):v\in[0,d],$
@@
4:  if $s_{t-1}=0$ or $b_{t-1}>S$ then $\triangleright$ refresh base drivers and storage manager driver
@@
For each driver $i\in\mathcal{B}\cup\mathcal{F}$, this is formalized as the definition of $\smash{\hat{x}^{(i)}_{t-1}}$ on Line 2 in Algorithm 1, where $q(x_{t-1})$ is the difference between the sum of all driver decisions and the global decision at time t1t-1. Intuitively, $q(x_{t-1})$ captures any “extra” purchasing from the previous time step that was not attributable to any currently active driver (e.g., due to base demand or inactive drivers). The pseudo-decision $\smash{\hat{x}_{t-1}^{(i)}}$ is used instead of the actual previous decision $\smash{x_{t-1}^{(i)}}$ when making the current decision $\smash{x_{t}^{(i)}}$. The same is applied to the delivery decisions of flexible drivers. Note that PAAD may generate as many as $\Omega(T)$ drivers on a given problem instance. The necessity of these many drivers is supported by our warmup result in § 3, which shows that a granular, per-demand accounting is likely necessary to achieve the best possible competitive ratio when future demands are unknown. However, PAAD’s driver refresh (Line 23) ensures that, in practice, there will likely be far fewer than $\Omega(T)$ drivers active at any given time. In particular, a driver is only active for a finite period—e.g., until flexible drivers’ deadlines pass, or until base drivers complete their purchasing. We discuss this point further and give runtime measurements in Appendix A.1.
4.2 Competitive Analysis
@@
$\alpha=\frac{\omega}{\Bigg[W\Bigg(-\frac{\left(P-(1+\varepsilon)p_{\min}\right)\exp\big(\frac{-\omega\nicefrac{{2\kappa}}{{T}}-p_{\min}c-P}{P+2\kappa}\big)}{P+2\kappa}\Bigg)+\frac{P+p_{\min}c-\omega\nicefrac{{2\kappa}}{{T}}}{P+2\kappa}\Bigg]},$ (2)
where $W(\cdot)$ is the Lambert W function Corless:96LambertW, $P=(1+c+\varepsilon)p_{\max}$, $\kappa=\gamma+\delta$ and $\omega=\frac{(1+c+\varepsilon)}{(1+\varepsilon)}$.
@@
$\alpha_{\texttt{T}}=\frac{\omega}{\Bigg[W\Bigg(-\frac{\left(P-(1+\varepsilon)p_{\min}\right)\exp\big(\frac{-\omega\nicefrac{{2\delta}}{{T}}-p_{\min}c-P}{P+2(\eta+\delta)}\big)}{P+2(\eta+\delta)}\Bigg)+\frac{P+p_{\min}c-\omega\nicefrac{{2\delta}}{{T}}}{P+2(\eta+\delta)}\Bigg]},$ (3)
@@ 4.3 Proof Overview
For an arbitrary OSDM instance, we then carefully lower bound the cost of OPT—compared to prior applications of the pseudo-cost minimization framework, the key technical challenge is that it is not sufficient to derive a single global lower bound on OPT’s cost (e.g., as in OCS Lechowicz:24), since there are many units of demand with differing constraints. The key idea of our analysis in Lemma E.3 is to partition the instance into active and inactive periods based on PAAD’s base driver state. This enables us to derive a lower bound on OPT’s cost that is a combination of the “local” best prices during each active period and the best price during all inactive periods, which is itself lower bounded. Next, we derive an upper bound on PAAD’s total cost by leveraging the definition of the pseudo-cost minimization to “charge” each driver’s decisions to their threshold functions (see Lemma E.4), while accounting for any worst-case purchasing, switching, and delivery costs that arise due to binding constraints for a particular unit of demand. To relate the cost of PAAD to that of OPT, we prove two lemmas that give a relation between each threshold function and its integral:
@@
The rest follows by contradiction—we show that if Lemma E.3 and E.4 hold, then if the threshold functions adhere to Lemmas 4.4 and 4.5, $\frac{\texttt{PAAD}(\mathcal{I})-p_{\max}\hat{s}}{\texttt{OPT}(\mathcal{I})}$ must be at most $\alpha$, completing the proof. $\square$
5 Fundamental Limits
In this section, given the results in § 4, we ask whether any algorithm can achieve a better competitive bound for OSDM than PAAD. In Theorems 5.1 and 5.3 below, we answer this question in the negative, showing that PAAD’s competitive ratio is the best achievable among all deterministic online algorithms for OSDM-S and OSDM-T, respectively. Along the way, we show results to contextualize the relative “hardness” of OSDM in subcases of interest, such as the case of only base demand (see Corollary 5.1) or the case of price-dependent increasing delivery cost (recall Def. 2.4).
@@ 5 Fundamental Limits
Proof Sketch of Theorem 5.1. In Def. F.1, we define a set of $x$-decreasing instances for OSDM-S (denoted by $\{\mathcal{I}_{x}\}_{x\in[p_{\min},p_{\max}]}$) where $x$ is the “best” (i.e., lowest) price in the instance. These instances consist of a single unit of flexible demand at time $t=1$ with a deadline of $\Delta_{1}=T$. An adversarial price sequence presents market prices that generally decrease over time, interrupted by “spikes” to the highest price $p_{\max}$ that force an online algorithm ALG to pay a high switching cost. This captures a trade-off between being eager or reluctant to accept a certain price $x$ – if ALG purchases too much at prices >x>x, it may miss better prices later on, but if it waits to purchase at prices <x<x, it may be forced to purchase and deliver at a high price $p_{\max}$ (e.g., if the sequence ends after $x$).
Under these special instances, the cost of any deterministic online algorithm ALG can be described by two (arbitrary) purchasing and delivery functions $h(x),z(x):[p_{\min},p_{\max}]\rightarrow[0,1]$ that describe the fraction of the demand purchased and delivered at prices $\geq x$. Since OPT is easy to describe on these instances (i.e., $\texttt{OPT}(\mathcal{I}_{x})=x+\frac{2\gamma}{T}+\varepsilon x+\frac{2\delta}{T})$, we show that the purchasing and delivery functions of an $\alpha^{\star}$-competitive ALG must satisfy a necessary condition that relates $h(x),z(x)$ to $\alpha^{\star}$ for all $x\in[p_{\min},p_{\max}]$. By applying Grönwall’s Inequality (Mitrinovic:91, Theorem 1, p.356), this gives the following condition that $\alpha^{\star}$ must satisfy at optimality: $\smash{1=\frac{\alpha^{\star}}{\omega}\ln\Big[\frac{P-(1+\varepsilon)p_{\min}}{P-\frac{\omega[P+2\kappa]}{\alpha^{\star}}+\frac{\omega 2\kappa}{T}+p_{\min}c}\Big]}$. Solving this transcendental for $\alpha^{\star}$ yields the result, completing the proof. $\square$
In the above result, we have shown that PAAD achieves the best possible competitive ratio for OSDM-S. The proof of Theorem 5.1 relies on a class of OSDM instances that consist purely of flexible demand. In the following corollary, we use a similar class of “hard instances” to argue that the competitive ratio improves if we focus on the case of OSDM-S instances with only base demand:
@@
$\alpha_{\texttt{B}}\geq\Bigg[W\Bigg(-\frac{\big((1+c)p_{\max}-p_{\min}\big)\exp\big(-\frac{P+2\delta+\frac{2\gamma}{T}}{P+2\kappa}\big)}{P+2\kappa}\Bigg)+\frac{P+2\delta+\frac{2\gamma}{T}}{P+2\kappa}\Bigg]^{-1}\kern-10.00002pt,$ (4)
@@
Delivery Cost Dynamics. In the above results and throughout the paper, we focus on the case of a price-dependent decreasing delivery cost. Amongst the class of monotone delivery costs defined in Def. 2.4, we claimed this is the most challenging case to consider in § 2. We now prove this claim by showing that OSDM with price-dependent increasing delivery cost is strictly easier:
@@
Amongst the class of monotone delivery costs defined in Def. 2.4, the best achievable competitive ratios for OSDM with an increasing delivery cost are strictly better than those for OSDM with a decreasing delivery cost (i.e., the cases considered in Theorem 5.1, 5.3 and Corollary 5.1).
We relegate the full proof of Theorem 5.2 to § F.3, which relies on lower bound “hard instance” constructions similar to those used in Theorem 5.1 and Corollary 5.1. The main insight is that since the case of price-dependent increasing delivery cost is strictly easier, our general results for the monotone delivery cost focus on the more challenging setting of decreasing delivery cost.
@@
6 Learning Data-Driven Decisions
@@
Recall that in § 2.1, we mentioned that algorithms designed using competitive analysis are often overly pessimistic, and can sometimes perform poorly compared to simple heuristics used in practice. However, these heuristics—as well as data-driven, machine-learned algorithms—can suffer a lack of robustness to nonstationarity or distribution shift. In this section, we propose and study a novel end-to-end learning methodology for OSDM that takes inspiration from recent literature on learning-augmented algorithms and learning-to-optimize (L2O) Chen:21:L20 to improve the average-case performance of online algorithms for OSDM while retaining worst-case robustness guarantees. We begin by reviewing some relevant preliminaries.
@@
Preliminaries & Challenges. In the literature on learning-augmented online algorithms, the competitive ratio (recall Def. 2.1) is usually replaced with two metrics: consistency and robustness Lykouris:18; Purohit:18; Mitzenmacher:22:ALPS. If ALG is a learning-augmented online algorithm provided with some advice ADV (e.g., a prediction about the instance), then ALG is said to be $c$-consistent if it is $c$-competitive when ADV is perfectly accurate, and $\rho$-robust if it is $\rho$-competitive, regardless of how bad (e.g., adversarial) ADV is.
There are two common models for advice. The first is decision advice, where ADV predicts the optimal decisions at each time step—this type of advice is frequently used for problems with multiple time steps and switching costs, such as metrical task systems Antoniadis:20MTS and the OCS problem discussed in § 3 Lechowicz:24 The second is input prediction advice, where ADV predicts a characteristic of the problem instance, such as the number of skiing days in a ski rental problem Purohit:18, or the best price in a sequence. This type of advice has been used for related problems such as online knapsack Daneshvaramoli:25 and one-way trading SunLee:21, which is itself closely related to OCS.
@@
Both models face limitations in the context of OSDM. Obtaining decision advice directly from an ML model can be difficult due to the generalization challenges posed by complex problem structure—existing methods instead predict input features (e.g., a price forecast in OCS), solve the corresponding “certainty equivalent” offline problem, and use that solution as advice Lechowicz:24. In OSDM, however, the joint uncertainty in prices and demand arrivals/deadlines makes this approach brittle to prediction error and impractical in practice. On the other hand, point predictions about instances (e.g., estimates of the best price) are straightforward to obtain, but these are too simplistic to provide useful insights for OSDM. For instance, even given a perfect prediction of the best (i.e., lowest) price, it is not possible to determine optimal purchasing and delivery schedules subject to the unknown demand arrivals/deadlines and switching/tracking costs.
@@
On top of these limitations to existing prediction models, there is a gap between theory and practice where performance metrics are concerned. In particular, while robustness is useful as a worst-case “guardrail” against, e.g., significantly out-of-distribution inputs, consistency is less meaningful—it is only defined for perfect predictions that are unrealistic in practice, and it is still a worst-case metric (i.e., it is defined as the competitive ratio given accurate predictions). While recent work has considered “smooth degradation” of algorithm performance to small prediction errors in online search problems Benomar:25; NEURIPS2024_11c6625b, it is not clear that these methods can be extended to the substantially more complex setting of OSDM. In practice, the primary objective is for an algorithm to perform well on average: if an algorithm performs very well in practice but has bad theoretical consistency, that is preferable to an algorithm that has good consistency in theory but poor real-world performance due to worst-case hedging. This motivates us to propose a new model that circumvents the challenges above while still providing the key benefits of “both worlds”—improved performance in practice while retaining worst-case guarantees.
6.1 The PALD Framework

### v2

@@ 4. A Unified Competitive Algorithm
4.1. The PAAD Algorithm
@@
Given a base driver with size denoted by $d$, the threshold function $\phi_{b}$ for OSDM-S is defined as follows, where $\alpha$ is the competitive ratio (see Thm. 4.3):
@@
$\psi(v)=p_{\max}(c+\varepsilon)+2\delta+\Big(\frac{p_{\max}(c+\varepsilon)+2\delta}{\alpha^{\prime}}-\Big(p_{\max}(c+\varepsilon)+\frac{2\delta}{T}\omega\Big)\Big)\exp\left(\frac{v}{\alpha^{\prime}d}\right):v\in[0,d].$
Figure 2. Diagram of the intuition behind the PAAD algorithm. Each driver accounts for a single unit of base or flexible demand, making local decisions based on its own state and role that are aggregated into a single global decision. Purchasing decisions $x_{t}$ govern how much to purchase at the current market price, and delivery decisions $z_{t}$ dictate how much demand is satisfied at the current time (subject to feasibility constraints).
@@
For each driver $i\in\mathcal{B}\cup\mathcal{F}$, this is formalized as the definition of $\smash{\hat{x}^{(i)}_{t-1}}$ on Line 2 in Algorithm 1, where $q(x_{t-1})$ is the difference between the sum of all driver decisions and the global decision at time t1t-1. Intuitively, $q(x_{t-1})$ captures any “extra” purchasing from the previous time step that was not attributable to any currently active driver (e.g., due to base demand or inactive drivers). Note that Line 3 (Algorithm 1) includes a $\gamma|x|$ term—this term accounts for the worst-case future cost of “switching off” whenever $x>0$ (e.g., when the sequence ends at time $T$, decisions $x_{T+1}=z_{T+1}=0$). By considering the pseudo-decision $\smash{\hat{x}_{t-1}^{(i)}}$ instead of the actual previous decision $\smash{x_{t-1}^{(i)}}$ when making the next decision, PAAD amortizes this “future cost” across multiple drivers, resulting in smoother global decisions. The same is applied to the delivery decisions of flexible drivers. Note that PAAD may generate as many as $\Omega(T)$ drivers on a given problem instance. The necessity of these many drivers is supported by our warmup result in § 3, which shows that a granular, per-demand accounting is likely necessary to achieve the best possible competitive ratio when future demands are unknown. However, PAAD’s driver refresh (Line 23) ensures that, in practice, there will likely be far fewer than $\Omega(T)$ drivers active at any given time. In particular, a driver is only active for a finite period—e.g., until flexible drivers’ deadlines pass, or until base drivers complete their purchasing. We discuss this point further and give runtime measurements in Appendix A.1.
@@
4:  if $s_{t-1}=0$ or $b_{t}>S$ then $\triangleright$ refresh base drivers and storage manager driver
@@
4.2. Competitive Analysis
@@
(3) $\alpha=\frac{\omega}{\Bigg[W\Bigg(-\frac{\left(P-(1+\varepsilon)p_{\min}\right)\exp\big(\frac{-\nicefrac{{\omega 2\kappa}}{{T}}-p_{\min}c-P}{P+2\kappa}\big)}{P+2\kappa}\Bigg)+\frac{P+p_{\min}c-\nicefrac{{\omega 2\kappa}}{{T}}}{P+2\kappa}\Bigg]},$
where $W(\cdot)$ is the Lambert W function (Corless et al., 1996), $P=(1+c+\varepsilon)p_{\max}$, $\kappa=\gamma+\delta$ and $\omega=\frac{(1+c+\varepsilon)}{(1+\varepsilon)}$.
@@
We remark on several regimes of interest for $\alpha$ where OSDM-S essentially “recovers” the optimal competitive ratio of other online problems. First, in a case with zero delivery cost (and zero delivery switching cost), letting $T\to\infty$, $\omega=1$, $\kappa=\beta$, $c=\epsilon=0$, and $P=U-2\beta$, [^3] In contrast to OSDM, the original OCS paper (Lechowicz et al., 2024) assumed that any necessary compulsory trading at the end of the time horizon could be accomplished without incurring additional switching cost, captured by the $2\beta$ in this term. $\alpha$ exactly reduces to the optimal competitive ratio for online conversion with switching costs, where $\beta$ is the switching coefficient (Lechowicz et al., 2024). Furthermore, letting $T\to\infty$, $\omega\to 1$, $\kappa=0$, $c=\epsilon=0$, and $P\to U$, $\alpha$ reduces to the optimal competitive ratio for the minimization version of one-way trading (Lorenz et al., 2008).
@@
(4) $\alpha_{\texttt{T}}=\frac{\omega}{\Bigg[W\Bigg(-\frac{\left(P-(1+\varepsilon)p_{\min}\right)\exp\big(\frac{-\nicefrac{{\omega 2\delta}}{{T}}-p_{\min}c-P}{P+2(\eta+\delta)}\big)}{P+2(\eta+\delta)}\Bigg)+\frac{P+p_{\min}c-\nicefrac{{\omega 2\delta}}{{T}}}{P+2(\eta+\delta)}\Bigg]},$
@@ 4.3. Proof Overview
For an arbitrary OSDM instance, we then carefully lower bound the cost of OPT—compared to prior applications of the pseudo-cost minimization framework, the key technical challenge is that it is not sufficient to derive a single global lower bound on OPT’s cost (e.g., as in OCS (Lechowicz et al., 2024)), since there are many units of demand with differing constraints. The key idea of our analysis in Lemma E.3 is to partition the instance into active and inactive periods based on PAAD’s base driver state. This enables us to derive a lower bound on OPT’s cost that is a combination of the “local” best prices during each active period and the best price during all inactive periods, which is itself lower bounded. Next, we derive an upper bound on PAAD’s total cost by leveraging the definition of the pseudo-cost minimization to “charge” each driver’s decisions to their threshold functions (see Lemma E.4), while accounting for any worst-case purchasing, switching, and delivery costs that arise due to binding constraints for a particular unit of demand. To relate the cost of PAAD to that of OPT, we prove two lemmas that give a relation between each threshold function and its integral:
@@
The rest follows by contradiction—we show that if Lemma E.3 and E.4 hold, then if the threshold functions adhere to Lemmas 4.5 and 4.6, $\frac{\texttt{PAAD}(\mathcal{I})-p_{\max}\hat{s}}{\texttt{OPT}(\mathcal{I})}$ must be at most $\alpha$, completing the proof. $\square$
5. Fundamental Limits
In this section, given the results in § 4, we ask whether any algorithm can achieve a better competitive bound for OSDM than PAAD. In Theorems 5.1 and 5.4 below, we answer this question in the negative, showing that PAAD’s competitive ratio is the best achievable among all deterministic online algorithms for OSDM-S and OSDM-T, respectively. Along the way, we show results to contextualize the relative “hardness” of OSDM in subcases of interest, such as the case of only base demand (see Corollary 5.2) or the case of price-dependent increasing delivery cost (recall Def. 2.4).
@@ 5. Fundamental Limits
Proof Sketch of Theorem 5.1. In Def. F.1, we define a set of $x$-decreasing instances for OSDM-S (denoted by $\{\mathcal{I}_{x}\}_{x\in[p_{\min},p_{\max}]}$) where $x$ is the “best” (i.e., lowest) price in the instance. These instances consist of a single unit of flexible demand at time $t=1$ with a deadline of $\Delta_{1}=T$. An adversarial price sequence presents market prices that generally decrease over time, interrupted by “spikes” to the highest price $p_{\max}$ that force an online algorithm ALG to pay a high switching cost. This captures a trade-off between being eager or reluctant to accept a certain price xx—if ALG purchases too much at prices >x>x, it may miss better prices later on, but if it waits to purchase at prices <x<x, it may be forced to purchase and deliver at a high price $p_{\max}$ (e.g., if the sequence ends after $x$).
Under these special instances, the cost of any deterministic online algorithm ALG can be described by two (arbitrary) purchasing and delivery functions $h(x),z(x):[p_{\min},p_{\max}]\rightarrow[0,1]$ that describe the fraction of the demand purchased and delivered at prices $\geq x$. Since OPT is easy to describe on these instances (i.e., $\texttt{OPT}(\mathcal{I}_{x})=x+\frac{2\gamma}{T}+\varepsilon x+\frac{2\delta}{T})$, we show that the purchasing and delivery functions of an $\alpha^{\star}$-competitive ALG must satisfy a necessary condition that relates $h(x),z(x)$ to $\alpha^{\star}$ for all $x\in[p_{\min},p_{\max}]$. By applying Grönwall’s Inequality (Mitrinovic et al., 1991, Theorem 1, p.356), this gives the following condition that $\alpha^{\star}$ must satisfy at optimality: $\smash{1=\frac{\alpha^{\star}}{\omega}\ln\Big[\frac{P-(1+\varepsilon)p_{\min}}{P-\frac{\omega[P+2\kappa]}{\alpha^{\star}}+\frac{\omega 2\kappa}{T}+p_{\min}c}\Big]}$. Solving this transcendental for $\alpha^{\star}$ yields the result, completing the proof. $\square$
In the above result, we have shown that PAAD achieves the best possible competitive ratio for OSDM-S. The proof of Theorem 5.1 relies on a class of OSDM instances that consist purely of flexible demand. In the following corollary, we use a similar class of “hard instances” to argue that the competitive ratio improves if we focus on the case of OSDM-S instances with only base demand—recall from § 2.1 that this case corresponds to a simplification of OSDM wherein the two decision variables of OSDM can be simplified to a single decision variable (i.e., $\{x_{t}-z_{t}\}_{t\in[T]}$).
@@
(5) $\alpha_{\texttt{B}}\geq\Bigg[W\Bigg(-\frac{\big((1+c)p_{\max}-p_{\min}\big)\exp\big(-\frac{P+2\delta+\frac{2\gamma}{T}}{P+2\kappa}\big)}{P+2\kappa}\Bigg)+\frac{P+2\delta+\frac{2\gamma}{T}}{P+2\kappa}\Bigg]^{-1}\kern-10.00002pt,$
@@
We remark that Theorem 5.1 and Corollary 5.2 do not rule out the possibility of randomization improving the achievable competitive ratio. While several works (El-Yaniv et al., 2001; Lorenz et al., 2008; Zhou et al., 2008) have shown that randomization does not improve the competitive ratio for continuous or fractional online allocation problems such as one-way trading and online knapsack, the smoothness terms in the objective of OSDM (and related problems such as OCS) may change this property.
Delivery Cost Dynamics. In the above results and throughout the paper, we focus on the case of a price-dependent delivery cost that is decreasing as a function of the storage state. Amongst the class of delivery costs defined in Def. 2.4, we claimed this is the most challenging case to consider in § 2. We now prove this claim by showing that OSDM with a price-dependent increasing delivery cost is strictly easier. As a corollary, any simpler special cases of Def. 2.4 (e.g., a price-dependent delivery cost that is constant in the storage state) are also easier in a competitive sense.
@@
Amongst the class of monotone delivery costs defined in Def. 2.4, the best achievable competitive ratios for OSDM with an increasing delivery cost are strictly better than those for OSDM with a decreasing delivery cost (i.e., the cases considered in Theorem 5.1, 5.4 and Corollary 5.2).
We relegate the full proof of Theorem 5.3 to § F.3, which relies on lower bound “hard instance” constructions similar to those used in Theorem 5.1 and Corollary 5.2. The main insight is that our general competitive results focus on the most challenging setting of a delivery cost that is price-dependent and decreasing as a function of the storage state—these results directly extend to the other “simpler” cases of delivery cost models.
@@
6. Learning Data-Driven Decisions
@@
In § 2.1, we mentioned that algorithms designed using competitive analysis are often overly pessimistic, and can sometimes perform poorly compared to simple heuristics used in practice. However, these heuristics—as well as data-driven, machine-learned algorithms—can suffer a lack of robustness to nonstationarity or distribution shift. In this section, we propose a novel paradigm that directly learns “the best OSDM algorithm” lying within a defined set of algorithms that satisfy a given (desired) worst-case performance guarantee, improving average-case performance while retaining robustness. To accomplish this, we show that the pseudo-cost design of PAAD provides a mechanism by which one can theoretically describe the general set of $\rho$-robust algorithms” (for any $\rho>\alpha$, see Def. 6.1). To realize this learning paradigm in practice, we show a end-to-end learning methodology (see § 6.2) that leverages recent techniques from the literature on learning-to-optimize (L2O) (Chen et al., 2021), such as differentiable optimization layers. We begin by reviewing some relevant preliminaries.
@@
Preliminaries & Challenges. In the literature on learning-augmented online algorithms, the competitive ratio (Def. 2.1) is replaced with two metrics: consistency and robustness (Lykouris and Vassilvtiskii, 2018; Kumar et al., 2018; Mitzenmacher and Vassilvitskii, 2022). If ALG is a learning-augmented online algorithm provided with some advice ADV (e.g., a prediction about the instance), then ALG is said to be $c$-consistent if it is $c$-competitive when ADV is perfectly accurate, and $\rho$-robust if it is $\rho$-competitive, regardless of how bad (e.g., adversarial) ADV is.
@@
There are two common models for advice. The first is decision advice, where ADV predicts the optimal decisions at each time step—this type of advice is frequently used for problems with multiple time steps and switching costs, such as metrical task systems (Antoniadis et al., 2020) and the OCS problem discussed in § 3 (Lechowicz et al., 2024) The second is input prediction advice, where ADV predicts a characteristic of the problem instance, such as the number of skiing days in a ski rental problem (Kumar et al., 2018). This type of advice has been used for related problems such as online knapsack (Daneshvaramoli et al., 2025) and one-way trading (Sun et al., 2021a), which is itself closely related to OCS. Both models face limitations in the context of OSDM. Obtaining decision advice directly from an ML model can be difficult due to the generalization challenges posed by complex problem structure—existing methods instead predict input features (e.g., a price forecast in OCS), solve the corresponding “certainty equivalent” offline problem, and use that solution as advice (Lechowicz et al., 2024). In OSDM, however, joint uncertainty in prices and demand arrivals/deadlines makes this approach brittle to prediction error and impractical in practice. On the other hand, point predictions about instances (e.g., estimates of the best price) are straightforward to obtain, but are too simplistic to provide useful insights for OSDM. For instance, even given a perfect prediction of the best (i.e., lowest) price, it is not possible to determine optimal purchasing and delivery schedules subject to the unknown demand arrivals/deadlines and switching/tracking costs.
@@
On top of these limitations to existing prediction models, there is a gap between theory and practice where metrics are concerned. In particular, while robustness is useful as a worst-case “guardrail” against, e.g., significantly out-of-distribution inputs, the consistency metric is less meaningful—it is only defined for perfect predictions that are unrealistic in practice, and it is still a worst-case metric (i.e., it is defined as the competitive ratio given accurate predictions). While recent work has considered “smooth degradation” of algorithm performance to small prediction errors in online search problems (Benomar et al., 2025; Elenter et al., 2024), it is not clear that these methods can be extended to the substantially more complex setting of OSDM. In practice, the actual objective is for an algorithm to perform well on average: if an algorithm performs very well in practice but has bad theoretical consistency, that is preferable to an algorithm that has good consistency in theory but poor real-world performance. This motivates us to propose a model that circumvents the challenges above while still providing the key benefits of “both worlds”—improved performance in practice with worst-case guarantees.
6.1. The PALD Framework